# EVCtrl: Efficient Control Adapter for Visual Generation

### v1

@@ 1 Introduction
Recent advances in the alignment of machine learning models for text-to-image [37, 32] and text-to-video synthesis [35, 36, 38, 34] have propelled a new wave of high-performance generators. Image models such as PixArt-deltadelta[6], Stable Diffusion 3[10], Flux[1], Kolors[17], and Hunyuan-DiT[20], together with video models including Sora[25], CogVideoX[59], HunyuanVideo[18], and Qihoo-T2X[50], now deliver stunning visual fidelity and open up a wide spectrum of downstream applications.
However, current ControlNet-based[63] controlled generation methods for DiT have two major drawbacks. First, as an ”auxiliary” neural network model structure, ControlNet typically directly replicates half of the network architecture, such as PixArt-$\delta$[6], where a large number of additional parameters significantly increase the inference burden. Similarly, although OminiControl[46] only adds a limited number of parameters, it doubles the number of tokens involved in the attention and linear layers, resulting in an almost 70% increase in overall computational complexity. Second, during the entire diffusion process, the correlation between different time steps and the feature changes of conditional tokens is often overlooked, leading to low computational efficiency. To address these issues caused by computational complexity, this paper first examines the redundancy of conditional information in time and space:
@@ 2.1 Controllable Diffusion Models
To achieve conditional control in pre-trained text-to-image diffusion models, there are two main methods for introducing controllable conditions into image [58, 56, 61, 65, 66] or video generation models [33, 31]: (1) training a new large model from scratch for multi-condition compliance [15] (2) freezing the large pretrained model and fine-tuning only a lightweight adapter. Recent studies have extended these ideas to diffusion transformers (DiTs) [45, 51, 4, 7, 5, 49, 12, 39, 60, 19, 43], exploiting the built-in multimodal attention (MM-attention)[41] to inject image or video conditions without further architectural change. T2I-Adapter[40] exemplifies the adapter route—it imposes minimal overhead yet yields weak control, limiting its use on complex prompts. In contrast, ControlNet[63] achieves stronger control effects by copying specific layers from pre-trained large models to guide image generation to align with control information, but it introduces a large number of additional parameters. Meanwhile, as the number of condition tokens increases, especially in controllable video generation, performing self-attention operations on the entire sequence of condition tokens leads to a significant increase in computational cost, imposing significant limitations on the controllable generation result rate from text to video.
@@ 2.2 Acceleration of Diffusion Models
Implementing low-latency and high-quality generation methods on diffusion models (DMs), including DiT, is an important research direction. Currently, there are mainly two types of acceleration methods for diffusion models: the first type is to reduce the number of sampling steps[24, 52, 26, 67, 8, 27, 44], and the second type is to accelerate the internal computation of diffusion models. Solutions to reduce computational complexity include model distillation and compression[55], token merging[3, 2, 11], token pruning[62], and layer-wise caching techniques[13, 54, 9, 22, 30, 29, 42]. However, layer-wise caching techniques use entire layers as the granularity and have difficulty perceiving the importance differences between tokens, resulting in key information being equally frozen or discarded. Toca[69] and Duca[70] instead use tokens as the minimum caching unit, accurately measure, and explicitly assign importance weights to each token with the help of a score map. In the subsequent computation stage, only high-weight tokens are refreshed according to a predetermined ratio, and the rest are directly reused from the cache, thus achieving lossless acceleration.
@@
Unfortunately, the existing caching schemes are not specifically designed to adapt to ControlNet, which replicates specific layers in DM, and have the following issues: First, the existing caching and editing schemes often rely on explicit read and write operations of attention maps[69] or KV matrices[28, 68], which are incompatible with the mainstream Transformer acceleration pipeline and instead introduce additional latency. Additionally, there are differences in the correlation and importance of control information among different transformer layers in the existing diffusion transformer controlnet[5]. Meanwhile, the control conditions themselves provide prior information on the spatial distribution that should be focused on, but current methods fail to utilize these differences and information.
@@ 4.1.1 Model Configurations
We conducted experiments with three commonly used DIT-based models equipped with ControlNet for various generative tasks, including Flux[57] for text-to-image generation, CogVideo[47], and Wan2.1[48] for text-to-video generation. The experiments were performed using an NVIDIA A800 80GB GPU. The CogVideo5B model is capable of generating 6-second videos with a resolution of 480p and a frame rate of 8fps. The Wan2.1-14B model can generate videos with a resolution of 720p or 480p, a frame rate of 16fps, and a duration of 5 seconds. Each model used its default sampling method: DPM-Solver++[27] with 20 steps for Flux, and DDIM[44] with 50 steps for both CogVideo and Wan2.1. For each model, we configure EVCtrl by setting the cache ratio $P$ according to the redundancy level of the conditional information, while also employing various average forced-activation periods $N$.
@@
We randomly selected 5,000 images and captions from the COCO-2017[21] dataset and extracted 100 videos and their descriptions from each of the 16 distinct categories of videos in the Vbench[16] dataset, totaling 1,600 videos. We evaluated the quality of the generated content using SSIM[53], LPIPS[64], PSNR[53], and FID[14]. Additionally, for text-to-video generation, we used video generation metrics within the Vbench framework as supplementary indicators.
@@ 4.2.1 Quantitative Results
As shown in Table 1, we benchmark EVCtrl against three state of the art training free acceleration baselines: Fora[42], Toca[69], and Taylorseer[23], on both the Flux-ControlNet[57] and CogVideo-ControlNet[47] pipelines. Quantitative results demonstrate that EVCtrl outperforms all other baselines across every quality metric (FID $\downarrow$, PSNR $\uparrow$, SSIM $\uparrow$, LPIPS $\downarrow$) while achieving equal or even higher acceleration ratios. Specifically, on the COCO2017 validation set at a $2.13\times$ speedup, EVCtrl achieves the lowest FID of 42.34 and maintains SSIM above 0.9; on the VBench validation set, it delivers a $2.16\times$ speedup with SSIM = 0.81, and all remaining metrics also surpass those of competing baselines at the same speedup, without any retraining or extra memory cost.
@@ 4.3.1 Effectiveness of Local Focused Caching(LFoC)
Table 3 shows that removing LFoC lowers SSIM from 0.81 to 0.66 and raises LPIPS from 0.15 to 0.27, indicating a decline in generation quality. Moreover, comparing LFoC-with-attention(WA) and LFoC-with-MLP(WM) reveals that updating either the attention layers alone or the MLP layers alone leaves the cached key control tokens under-utilized. Only when LFoC refreshes both the attention and MLP layers simultaneously can the re-injected tokens fully unleash their expressive power, allowing the model to recover its original quality metrics. The visual results in Figure 5 demonstrate that after removing LFoC the model struggles to control fine details, illustrating that LFoC better exploits spatially local, fine-grained prior knowledge and achieves superior controllability in controlled generation tasks compared with the aggressive cache approach and the approach [69][70] that only updates the MLP layers.

### v2

@@ 1 Introduction
Recent advances in the alignment of machine learning models for text-to-image and text-to-video synthesis have propelled a new wave of high-performance generators. Image models such as PixArt-deltadelta[chen2024pixartdelta], Stable Diffusion 3[esser2024scaling], Flux[flux2024], Kolors[kolors], and Hunyuan-DiT[li2024hunyuandit], together with video models including Sora[liu2024sorareviewbackgroundtechnology], CogVideoX[yang2024cogvideox], HunyuanVideo[kong2024hunyuanvideo], and Qihoo-T2X[wang2024qihoot2xefficientproxytokenizeddiffusion], now deliver stunning visual fidelity and open up a wide spectrum of downstream applications.
However, current ControlNet-based[zhang2023adding] controlled generation methods for DiT have two major drawbacks. First, as an ”auxiliary” neural network model structure, ControlNet typically directly replicates half of the network architecture, such as PixArt-$\delta$[chen2024pixartdelta], where a large number of additional parameters significantly increase the inference burden. Similarly, although OminiControl[tan2025ominicontrol] only adds a limited number of parameters, it doubles the number of tokens involved in the attention and linear layers, resulting in an almost 70% increase in overall computational complexity. Second, during the entire diffusion process, the correlation between different time steps and the feature changes of conditional tokens is often overlooked, leading to low computational efficiency. To address these issues caused by computational complexity, this paper first examines the redundancy of conditional information in time and space:
@@ 2.1 Controllable Diffusion Models
To achieve conditional control in pre-trained text-to-image diffusion models, there are two main methods for introducing controllable conditions into image or video generation models: (1) training a new large model from scratch for multi-condition compliance [huang2023composer] (2) freezing the large pretrained model and fine-tuning only a lightweight adapter. Recent studies have extended these ideas to diffusion transformers (DiTs) [tan2024ominicontrol, cai2025diffusion, cao2025relactrl, mao2025ace++, she2025customvideox], exploiting the built-in multimodal attention (MM-attention)[pan2020multi] to inject image or video conditions without further architectural change. T2I-Adapter[mou2024t2i] exemplifies the adapter route—it imposes minimal overhead yet yields weak control, limiting its use on complex prompts. In contrast, ControlNet[zhang2023adding] achieves stronger control effects by copying specific layers from pre-trained large models to guide image generation to align with control information, but it introduces a large number of additional parameters. Meanwhile, as the number of condition tokens increases, especially in controllable video generation, performing self-attention operations on the entire sequence of condition tokens leads to a significant increase in computational cost, imposing significant limitations on the controllable generation result rate from text to video.
@@ 2.2 Acceleration of Diffusion Models
Implementing low-latency and high-quality generation methods on diffusion models (DMs), including DiT, is an important research direction. Currently, there are mainly two types of acceleration methods for diffusion models: the first type is to reduce the number of sampling steps[liu2022flow, lu2022dpm, lu2025dpm, song2020denoising], and the second type is to accelerate the internal computation of diffusion models. Solutions to reduce computational complexity include model distillation and compression[xie2024sana], token merging[bolya2022token, bolya2023token, feng2025dit4edit], token pruning[zhang2025training], and layer-wise caching techniques[gao2024dual, liu2024faster, ma2024deepcache, ma2024learning, selvaraju2024fora]. However, layer-wise caching techniques use entire layers as the granularity and have difficulty perceiving the importance differences between tokens, resulting in key information being equally frozen or discarded. Toca[zou2024accelerating] and Duca[zou2024accelerating2] instead use tokens as the minimum caching unit, accurately measure, and explicitly assign importance weights to each token with the help of a score map. In the subsequent computation stage, only high-weight tokens are refreshed according to a predetermined ratio, and the rest are directly reused from the cache, thus achieving lossless acceleration.
@@
Unfortunately, the existing caching schemes are not specifically designed to adapt to ControlNet, which replicates specific layers in DM, and have the following issues: First, the existing caching and editing schemes often rely on explicit read and write operations of attention maps[zou2024accelerating] or KV matrices[lu2023tf, zhu2025kv], which are incompatible with the mainstream Transformer acceleration pipeline and instead introduce additional latency. Additionally, there are differences in the correlation and importance of control information among different transformer layers in the existing diffusion transformer controlnet[cao2025relactrl]. Meanwhile, the control conditions themselves provide prior information on the spatial distribution that should be focused on, but current methods fail to utilize these differences and information.
@@ 4.1.1 Model Configurations
We conducted experiments with three commonly used DIT-based models equipped with ControlNet for various generative tasks, including Flux[flux_controlnet] for text-to-image generation, CogVideo[cogvideo_controlnet], and Wan2.1[wan2025wan] for text-to-video generation. The experiments were performed using an NVIDIA A800 80GB GPU. The CogVideo5B model is capable of generating 6-second videos with a resolution of 480p and a frame rate of 8fps. The Wan2.1-14B model can generate videos with a resolution of 720p or 480p, a frame rate of 16fps, and a duration of 5 seconds. Each model used its default sampling method: DPM-Solver++[lu2025dpm] with 20 steps for Flux, and DDIM[song2020denoising] with 50 steps for both CogVideo and Wan2.1. For each model, we configure EVCtrl by setting the cache ratio $P$ according to the redundancy level of the conditional information, while also employing various average forced-activation periods $N$.
@@
We randomly selected 5,000 images and captions from the COCO-2017[lin2014microsoft] dataset and extracted 100 videos and their descriptions from each of the 16 distinct categories of videos in the Vbench[huang2024vbench] dataset, totaling 1,600 videos. We evaluated the quality of the generated content using SSIM[wang2004image], LPIPS[zhang2018unreasonable], PSNR[wang2004image], and FID[heusel2017gans]. Additionally, for text-to-video generation, we used video generation metrics within the Vbench framework as supplementary indicators.
@@ 4.2.1 Quantitative Results
As shown in Table 1, we benchmark EVCtrl against three state of the art training free acceleration baselines: Fora[selvaraju2024fora], Toca[zou2024accelerating], and Taylorseer[TaylorSeer2025], on both the Flux-ControlNet[flux_controlnet] and CogVideo-ControlNet[cogvideo_controlnet] pipelines. Quantitative results demonstrate that EVCtrl outperforms all other baselines across every quality metric (FID $\downarrow$, PSNR $\uparrow$, SSIM $\uparrow$, LPIPS $\downarrow$) while achieving equal or even higher acceleration ratios. Specifically, on the COCO2017 validation set at a $2.13\times$ speedup, EVCtrl achieves the lowest FID of 42.34 and maintains SSIM above 0.9; on the VBench validation set, it delivers a $2.16\times$ speedup with SSIM = 0.81, and all remaining metrics also surpass those of competing baselines at the same speedup, without any retraining or extra memory cost.
@@ 4.3.1 Effectiveness of Local Focused Caching(LFoC)
Table 3 shows that removing LFoC lowers SSIM from 0.81 to 0.66 and raises LPIPS from 0.15 to 0.27, indicating a decline in generation quality. Moreover, comparing LFoC-with-attention(WA) and LFoC-with-MLP(WM) reveals that updating either the attention layers alone or the MLP layers alone leaves the cached key control tokens under-utilized. Only when LFoC refreshes both the attention and MLP layers simultaneously can the re-injected tokens fully unleash their expressive power, allowing the model to recover its original quality metrics. The visual results in Figure 5 demonstrate that after removing LFoC the model struggles to control fine details, illustrating that LFoC better exploits spatially local, fine-grained prior knowledge and achieves superior controllability in controlled generation tasks compared with the aggressive cache approach and the approach [zou2024accelerating][zou2024accelerating2] that only updates the MLP layers.
@@ 5 Conclusion
Appendix A Prompt Details
In Figure 1, the prompts for the videos and images are arranged top-to-bottom, left-to-right as follows:(1)“A cute panda sits on a tree branch” (2)“A young man is joyfully dancing in his room”(3)“Bright high-tech factory with automated electric-car production line”(4)“A freight truck cruises along an international highway”(5)“Scandinavian cozy interior, pale ash wood and warm candle-lit cream tones, ultra-HD, cinematic mood”(6)“Moonlit galleon sailing beneath bright moonlight”(7)“A white sedan cruises along a cliff-hugging coastal road”(8)“Masterpiece, best_quality, absurdres, conceptual art, landscape, mountains, fantasy, waterfalls, green, lush, valley, dramatic, artistic, illustration, nature, rocks, clouds, mystical, vibrant, aerial view, beach, forest, ocean, coastline, tropical, sand, waves, blue water, green trees, drone shot”(9)“Futuristic 500m shopping tower, glass curves, warm lights, night city”(10)“On the table sit chocolate, butter, bread, eggs, and a jug of milk”
In Figure 2, the video prompt is: “A majestic three-masted sailing ship is sailing across the vast ocean under a bright and sunny sky.”
In Figure 5, the video prompt is:“The Alhambra in daylight.”
Appendix B Related Works
In completing this work, we have systematically drawn upon a curated selection of distinguished studies from the image and video generation literature [yan2025eedit, xiong2025enhancing, ma2025followyourclick, yuluo2025gr, zhang2025magiccolor, ma2023magicstick, zhu2024instantswap, ma2025controllable, long2025follow, ma2024followpose, li2025skipvar, ma2022visual, wan2024unipaint, feng2025follow, ma2025followcreation, wang2024cove, chen2024follow, ma2025followyourmotion, wang2024taming, zhu2025multibooth, chen2023attentive, ma2024followyouremoji, xiao2023semanticac, chen2024m].These prior contributions constitute an integral component of the intellectual scaffolding underlying the present study.
Appendix C Experimental Details
C.1 PowerPoint for Video Presentation
To further demonstrate the superiority of EVCtrl in video generation, we have prepared an anonymous PowerPoint presentation. Please refer to the PowerPoint file in the Supplementary Materials for more detailed viewing.